OpenStudy (anonymous):

PLEASEEE HELPP IM GIVE MEDALSSS!!!! Evaluate the following expression using the values given. Find 2x2 - 2z4 + y2 - x2 + z4 if x = -4, y = 3, and z = 2.

OpenStudy (anonymous):

@zaibali.qasmi

OpenStudy (anonymous):

1st simplfy the relation..by adding same terms as.. 3x2-z4+y2 now just plug in the vales... 3(-4)^2-(2)^4+(3)^2 =3(16)-16+9=48-16+9=41

OpenStudy (anonymous):

thanks babes can u help me on this one ?Evaluate the following expression using the values given: Find 3x - y - 3z if x = -2, y = 1, and z = -2

OpenStudy (anonymous):

ummm..why not sweet baby... sam plug in values and get result..do it and tell me what you get..??